# CNN on ‘Top’: In Search of Scalable & Lightweight Image-based Jet Taggers

### v1

@@ 2.3 Data: Jet-Level Information
We have combined some global features of the jet to ascertain how much they improve the tagging efficiency of the CNNs. The features used in this work are inspired by ref.s ATL-PHYS-PUB-2021-028 ; Bhattacherjee:2022gjq . The original dataset provides the four-momentum of up to 200 constituents per jet, from which all global features can be computed. For this purpose, we employ the fastjet package along with EnergyCorrelator contrib from fastjet-contrib-1.3.2.
In addition to the global four-momentum of the jet, our feature set includes the N-subjettiness variables Thaler:2010tr , as well as several series of observables constructed from Energy Correlation Functions Larkoski:2013eya and generalized correlator functions Moult:2016cvt . The complete list of variables is as follows:
@@
  • $D$-series [^4] The parameter choices for the D-series features follow Larkoski:2014gra ; Larkoski:2015kga ; Larkoski:2014zma .: $D_{2}^{\alpha,\beta}$ for $\alpha,\beta=1.0,\,2.0$, and the four $D_{3}$ variables defined in Bhattacherjee:2022gjq
@@
  • L-series: $L_{2},L_{3}$ as defined in ATL-PHYS-PUB-2021-028
Certain global features, such as $U_{3}^{\beta}$ and $M_{3}^{\beta}$, require at least 4 constituents for their computation. As mentioned in the previous sub-section, we discard all jets with $N_{\text{constituents}}<4$. For complete definitions and a detailed description of these variables, we refer the reader to ref.s Bhattacherjee:2022gjq ; ATL-PHYS-PUB-2021-028
@@ 2.4 Model Architecture
Historically, one of the most successful convolutional neural networks (CNNs), LeNet lenetPaper , is also the oldest and one of the simplest. LeNet-5, with alternating convolutional and pooling layers, has been the dependable standard for smaller images (see the left side of Fig. (2) for its architecture). Over the decades following LeNet, CNNs continued to scale up to classify images at increasingly higher resolutions. AlexNet AlexNetPaper , Inception nets InceptionV1 ; InceptionV2V3 , ResNets ResNet1 and their wider versions wideResNet , DenseNets DenseNet , and MobileNets MobileNetV1 ; MobileNetV2 ; MobileNetV3 ; MobileNetV4 are only a few examples of these. To go deeper and wider, these newer architectures used varied techniques, such as ‘residual’ connections to feed information from earlier layers to the latter parts of the network, point-wise $(1\times 1)$ and depthwise convolutions for dimension reduction to remove computational bottlenecks, batch normalization, and mixtures of these methods InceptionV4 ; ResNext . Targeting general-purpose image classification, these models were optimized for the ImageNet dataset ImageNet2019 ; ImageNet2021 and its equivalents; thus, computationally, they are quite expensive.
@@
Network Input Image No. of Inference Time Accuracy AUC Bkg. Rej at TPR
Type Array Size Resolution Param.s /100 Jets ($ms$) (%) (%) $50\%$ $30\%$
LeNet $3\times 28\times 28$ $32\to 28$ $428K$ $5.5(7)$ $92.52$ $97.81$ $160(2)$ $525(23)$
(Images) $3\times 32\times 32$ $653K$ $7.3(8)$ $92.54$ $97.80$ $162(2)$ $536(14)$
$3\times 40\times 40$ $1253K$ $6.5(20)$ $92.56$ $97.82$ $169(3)$ $568(31)$
$3\times 32\times 32$ $40\to 32$ $653K$ $6.7(6)$ $92.55$ $97.82$ $165(4)$ $516(20)$
$3\times 40\times 40$ $1253K$ $6.5(7)$ $92.51$ $97.75$ $155(2)$ $476(16)$
$3\times 28\times 28$ $32\to 28$ $485K$ $6.4(6)$ $93.12$ $98.11$ $225(4)$ $774(33)$
$3\times 32\times 32$ $710K$ $7.5(5)$ $93.10$ $98.11$ $222(6)$ $777(35)$
(Images + $3\times 40\times 40$ $1310K$ $7.3(7)$ $93.08$ $98.11$ $222(3)$ $795(23)$
Global Feat.s) $3\times 32\times 32$ $40\to 32$ $710K$ $7.0(7)$ $93.17$ $98.14$ $233(2)$ $793(19)$
$3\times 40\times 40$ $1310K$ $7.7(8)$ $\mathbf{93.20}$ $\mathbf{98.15}$ $237(3)$ $\mathbf{814(24)}$
EffNet-S $3\times 40\times 40$ $32\to 28$ $162K$ $7.7(7)$ $92.59$ $97.85$ $174(5)$ $\mathbf{602(14)}$
(Images) $3\times 64\times 64$ $305K$ $10.4(6)$ $\mathbf{92.70}$ $\mathbf{97.91}$ $170(13)$ $577(67)$
$3\times 40\times 40$ $40\to 32$ $162K$ $8.5(6)$ $91.97$ $97.51$ $129(9)$ $415(40)$
(Images + $3\times 40\times 40$ $32\to 28$ $208K$ $8.4(6)$ $93.12$ $98.13$ $233(6)$ $\mathbf{864(44)}$
Global Feat.s) $3\times 64\times 64$ $365K$ $12.0(11)$ $\mathbf{93.13}$ $98.15$ $226(8)$ $817(38)$
DeepTop Macaluso:2018tck ; Kasieczka:2017nvn ; Kasieczka:2019dbj $1\times 37\times 37$ $60\to 40\to 37$ $610K$ - $93.00$ $98.10$ - $914(14)$
ResNeXt-50 ResNext
Ref. Kasieczka:2019dbj $1\times 64\times 64$ $96\to 64$ $1460K$ - $93.60$ $98.40$ - $1122(47)$
Ref. Qu:2019gqs $1\times 64\times 64$ $96\to 64$ $1460K$ 22 $93.60$ $98.37$ $302(5)$ $1147(58)$
@@
Table 2: Single number performance metrics for ML models performed on the total test dataset. From the second to the seventh column, we respectively report the input array sizes, the resolution of the generated jet images ($A\to B$ means the original image was $A\times A$, and it was cropped to $B\times B$), total number of parameters in the network, inference time for 100 jets in milliseconds, overall accuracy, and area under the ROC curve (AUC). The last two columns list the QCD background rejection at top tagging efficiencies of $50\%$ and $30\%$ respectively.
@@
Though scaling CNNs according to image resolution has been attempted before long2015fully , EfficientNets EffNetV1 ; EffNetV2 were the first to solve the scaling problem by using ‘compound scaling’, connecting network depth and width to image resolution. Given a baseline network (found from a neural architecture search, or NAS), these quantities scale as:
@@
where $\alpha\geq 1$, $\beta\geq 1$, $\gamma\geq 1$. Here, increasing $\phi$, the scaling parameter, increases the complexity of the network in a compounded manner. Noting that doubling the net-depth doubles the computational cost (floating point operations per second, or FLOPS), but doubling the other two quadruples the FLOPS, these also follow an empirical constraint: $\alpha\cdot\beta^{2}\cdot\gamma^{2}\simeq 2$, so that increasing $\phi$ by some $\phi_{0}$ increases the FLOPS by $2^{\phi_{0}}$. The authors of the first EfficientNet paper performed a grid search on $\alpha$, $\beta$, and $\gamma$ and obtained the values $1.2$, $1.1$, and $1.15$, respectively. Starting with $\phi=0$ and increasing $\phi$ by 1 each time, they obtained eight networks ($B_{0}$ to $B_{7}$) with increasing complexity. The base network ($B_{0}$) works on arrays of dimension $224\times 224$, while the largest of these models ($B_{7}$) takes in arrays of dimension $600\times 600$. There are seven main building blocks of an EfficientNet, each containing multiple parts. Building upon other successful lightweight models MobileNetV2 ; MNASnet , EfficientNets replace a fully convolutional operation with a depthwise convolution (a single filter per channel) and a pointwise ($1\times 1$) convolution (see the blown up schematic of Block 2A in Fig. (2) ). This significantly reduces the computational cost while maintaining similar performance. When stacked together, the blocks mentioned above also perform a function known as inverted bottlenecking, embedding the input manifold into a lower dimensional manifold. Each of these blocks also contains a squeeze and expansion block squeeze , with a residual connection maintaining the information flow, as well as a combination of Swish, RELU, and Sigmoid activation functions. More details can be found in the original paper EffNetV1 . While the width coefficient scales up the channel dimension of each block, the depth coefficient determines the number of times each type of block is repeated, making the network deeper. These networks also scale up the dropout probability to regularize the larger models. All versions of EfficientNet outperform other CNNs that are much larger in size (in terms of accuracy) while being faster during inference ($B_{0}$, with 5.3M5.3~M parameters, achieves a top-5 accuracy of $93.3\%$ for ImageNet, compared to the widely used ResNet-50 (26M26~M parameters)—which has a top-5 accuracy of $93.0\%$; the FLOPS for ResNet-50 is $11$ times that of EfficientNet-$B_{0}$).
@@
Figure 3: ROC curve (signal efficiency vs. background rejection) for some of the networks (with some metric highest among others; table (2) has them in bold-face). The solid lines are for models working on both images and global features, dashed ones correspond to those with only images as input.
To circumvent this without changing the hyperparameters of the original EfficientNet architecture (and using the prescribed values of $\alpha$, $\beta$, and $\gamma$), we use negative $\phi$ values to scale down the input resolution. The smallest resolution thus achieved is $36\times 36$. Any smaller, and the bottlenecks within the network would not hold without modification. We choose the starting resolution to be $40\times 40$ (in accordance with the resolution of the simulated calorimeter), corresponding to $\phi\simeq-12.33$. A larger version was also created with $64\times 64$ inputs, for $\phi\simeq-8.96$ [^6] Similar efforts can be found here: https://github.com/qubvel/efficientnet/issues/26. The other important parameter to set is the dropout probability, which is gradually increased from $B_{0}$ to $B_{7}$, as more regularization is needed for larger networks. We linearly extrapolated these values using the width and depth coefficients and used the corresponding values for our networks. Evidently, the much smaller size of our networks resulted in significantly less dropout regularization. Altogether, we will denote this smaller architecture as EfficientNet-Small (EffNet-S) for this work. For these values of $\phi$, every block appears only once in the smaller EffNet-S structures. We have thus indexed the blocks ‘$i$-A’ ($i=1,\,2,\,\ldots$), to keep the possibility of larger networks open [^7] This raises an obvious question: as these CNNs were fine-tuned to detect specific shapes (with detail) from larger images, unlike in our case—where the jet-images are smaller and are mere representations with a much larger importance to finer structures—could it be possible to systematically search for and create versions of scalable networks for this specific task, albeit smaller? Though intriguing, this question is outside the scope of the present work. However, glimpses of this possibility may be perceived from the present analysis.. Table (1) lists the input dimensions for this architecture for two different input resolutions: $40\times 40$ and $64\times 64$.

### v2

@@ 2.3 Data: Jet-Level Information
We have combined some global features of the jet to ascertain how much they improve the tagging efficiency of the CNNs. The features used in this work are inspired by ref.s 38; B. Bhattacherjee, C. Bose, A. Chakraborty, and R. Sengupta (2024). The original dataset provides the four-momentum of up to 200 constituents per jet, from which all global features can be computed. For this purpose, we employ the fastjet package along with EnergyCorrelator contrib from fastjet-contrib-1.3.2.
In addition to the global four-momentum of the jet, our feature set includes the N-subjettiness variables Thaler and Van Tilburg (2011), as well as several series of observables constructed from Energy Correlation Functions Larkoski et al. (2013) and generalized correlator functions Moult et al. (2016). The complete list of variables is as follows:
@@
  • $D$-series [^4] The parameter choices for the D-series features follow Larkoski et al. (2014, 2016, 2015).: $D_{2}^{\alpha,\beta}$ for $\alpha,\beta=1.0,\,2.0$, and the four $D_{3}$ variables defined in Bhattacherjee et al. (2024)
@@
  • L-series: $L_{2},L_{3}$ as defined in 38
Certain global features, such as $U_{3}^{\beta}$ and $M_{3}^{\beta}$, require at least 4 constituents for their computation. As mentioned in the previous sub-section, we discard all jets with $N_{\text{constituents}}<4$. For complete definitions and a detailed description of these variables, we refer the reader to ref.s B. Bhattacherjee, C. Bose, A. Chakraborty, and R. Sengupta (2024); 38
@@ 2.4 Model Architecture
Historically, one of the most successful convolutional neural networks (CNNs), LeNet Lecun et al. (1998), is also the oldest and one of the simplest. LeNet-5, with alternating convolutional and pooling layers, has been the dependable standard for smaller images (see the left side of Fig. (2) for its architecture). Over the decades following LeNet, CNNs continued to scale up to classify images at increasingly higher resolutions. AlexNet Krizhevsky et al. (2012), Inception nets Szegedy et al. (2015, 2016), ResNets He et al. (2016) and their wider versions Zagoruyko and Komodakis (2016), DenseNets Huang et al. (2017), and MobileNets Howard et al. (2017); Sandler et al. (2018); Howard et al. (2019); Qin et al. (2024) are only a few examples of these. To go deeper and wider, these newer architectures used varied techniques, such as ‘residual’ connections to feed information from earlier layers to the latter parts of the network, point-wise $(1\times 1)$ and depthwise convolutions for dimension reduction to remove computational bottlenecks, batch normalization, and mixtures of these methods Szegedy et al. (2017); Xie et al. (2017). Targeting general-purpose image classification, these models were optimized for the ImageNet dataset Yang et al. (2020, 2022) and its equivalents; thus, computationally, they are quite expensive.
@@
Network Input Img. Res. No. of Inference Accuracy AUC Bkg. Rejection
Type Array ($Org\to$ Param.s Time ($ms$) (%) at TPR
($3\times n\times n$) CrpCrp) Param.s /100 Jets Avg. Best Avg. Best $50\%$ $30\%$
$28$ $35\to 28$ $428K$ $5.7(8)$ $92.42(3)$ $92.48$ $97.74(1)$ $97.76$ $156(2)$ $480(18)$
LeNet $32$ $653K$ $6.1(9)$ $92.40(2)$ $92.43$ $97.74(1)$ $97.74$ $154(3)$ $483(13)$
(Images) $40$ $1253K$ $6.4(8)$ $92.45(3)$ $\mathbf{92.50}$ $97.76(2)$ $\mathbf{97.78}$ $\mathbf{161(4)}$ $\mathbf{514(23)}$
$32$ $40\to 32$ $653K$ $11.9(10)$ $92.41(4)$ $92.48$ $97.75(1)$ $97.76$ $161(3)$ $512(25)$
$40$ $1253K$ $13.4(18)$ $92.42(5)$ $92.48$ $97.74(1)$ $97.74$ $154(3)$ $476(11)$
$28$ $35\to 28$ $485K$ $5.7(8)$ $93.33(2)$ $93.37$ $98.24(0)$ $98.24$ $258(5)$ $916(25)$
(Images $32$ $710K$ $6.1(8)$ $93.34(2)$ $\mathbf{93.38}$ $98.24(1)$ $\mathbf{98.26}$ $257(3)$ $942(31)$
+ $40$ $1310K$ $6.6(9)$ $93.31(3)$ $93.36$ $98.23(1)$ $98.25$ $255(5)$ $909(34)$
Global $32$ $40\to 32$ $710K$ $14.5(16)$ $93.33(2)$ $93.36$ $98.24(0)$ $98.24$ $255(2)$ $\mathbf{951(43)}$
Feat.s) $40$ $1310K$ $14.8(7)$ $93.34(2)$ $93.37$ $98.24(1)$ $98.25$ $\mathbf{263(3)}$ $906(32)$
$36$ (S0) $35\to 28$ $158K$ $7.3(9)$ $92.31(11)$ $92.50$ $97.73(4)$ $97.80$ $160(5)$ $536(24)$
EffNet-S $40$ (S1) $184K$ $8.0(7)$ $92.35(8)$ $92.46$ $97.75(3)$ $97.76$ $163(3)$ $552(18)$
$48$ (S2) $215K$ $8.7(8)$ $92.39(9)$ $92.53$ $97.76(4)$ $97.80$ $\mathbf{167(5)}$ $569(30)$
$56$ (S3) $258K$ $9.6(8)$ $92.58(1)$ $92.72$ $97.85(3)$ $97.90$ $171(6)$ $\mathbf{585(32)}$
(Images) $64$ (S4) $305K$ $10.6(8)$ $92.20(40)$ $\mathbf{92.75}$ $97.74(11)$ $\mathbf{97.92}$ $156(16)$ $499(78)$
$36$ (S0) $40\to 32$ $158K$ $13.5(10)$ $91.40(60)$ $91.91$ $97.37(13)$ $97.50$ $127(11)$ $431(49)$
$40$ (S1) $184K$ $13.9(10)$ $90.5(12)$ $91.80$ $97.22(26)$ $97.39$ $112(15)$ $361(57)$
$48$ (S2) $215K$ $14.8(10)$ $91.4(10)$ $92.05$ $97.35(23)$ $97.56$ $126(18)$ $407(64)$
$56$ (S3) $258K$ $15.8(7)$ $92.0(4)$ $92.31$ $97.60(10)$ $97.68$ $143(9)$ $460(45)$
$64$ (S4) $305K$ $16.5(9)$ $91.9(6)$ $92.42$ $97.52(18)$ $97.74$ $137(17)$ $434(63)$
$36$ (S0) $35\to 28$ $242K$ $8.2(9)$ $93.28(2)$ $93.32$ $98.21(1)$ $\mathbf{98.23}$ $242(5)$ $890(34)$
(Img.s + $40$ (S1) $276K$ $7.9(9)$ $93.26(2)$ $93.30$ $98.21(1)$ $98.22$ $241(7)$ $859(34)$
Glob. Feat.s) $48$ (S2) $315K$ $8.4(9)$ $93.20(2)$ $93.30$ $98.18(2)$ $98.20$ $241(10)$ $839(51)$
$56$ (S3) $368K$ $10.0(10)$ $93.31(5)$ $\mathbf{93.38}$ $98.21(1)$ $98.22$ $232(5)$ $817(26)$
$64$ (S4) $422K$ $11.4(9)$ $93.15(5)$ $93.24$ $98.11(3)$ $98.15$ $208(8)$ $707(48)$
$36$ (S0) $40\to 32$ $242K$ $13.8(8)$ $93.24(8)$ $93.31$ $98.20(4)$ $\mathbf{98.23}$ $244(13)$ $\mathbf{897(74)}$
$40$ (S1) $276K$ $13.6(12)$ $93.26(2)$ $93.31$ $98.20(1)$ $98.21$ $\mathbf{247(6)}$ $868(25)$
$48$ (S2) $315K$ $14.4(9)$ $93.20(3)$ $93.26$ $98.17(2)$ $98.20$ $235(6)$ $864(30)$
$56$ (S3) $368K$ $15.7(9)$ $93.22(9)$ $93.35$ $98.18(3)$ $98.23$ $237(9)$ $829(44)$
$64$ (S4) $422K$ $17.0(10)$ $93.21(5)$ $93.27$ $98.15(3)$ $98.18$ $225(13)$ $778(71)$
DeepTop $1\times$ $60\to 40$
Macaluso and Shih (2018); Kasieczka et al. (2017); Butter and others (2019) $\times 37 37\to 37$ $610K$ $93.00$ $98.10$ $914(14)$
ResNeXt-50 Xie et al. (2017)
Ref. Butter and others (2019) $1\times$
$64\times 64$ $96\to 64$ $1460K$ $93.60$ $98.40$ $1122(47)$
Ref. Qu and Gouskos (2020) $1\times$
$64\times 64$ $96\to 64$ $1460K$ 22 $93.60$ $98.37$ $302(5)$ $1147(58)$
@@
Table 2: Single number performance metrics for ML models performed on the total test dataset. From the second to the seventh column, we respectively report the input array sizes, the resolution of the generated jet images ($A\to B$ means the original image was $A\times A$, and it was cropped to $B\times B$), total number of parameters in the network, inference time for 100 jets in milliseconds, overall accuracy, and area under the ROC curve (AUC). The last two columns list the QCD background rejection at top tagging efficiencies of $50\%$ and $30\%$, respectively. The values in bold-face represent the best metrics for a specific type of Network.
@@
Though scaling CNNs according to image resolution has been attempted before Long et al. (2015), EfficientNets Tan and Le (2019, 2021) were the first to solve the scaling problem by using ‘compound scaling’, connecting network depth and width to image resolution. Given a baseline network (found from a neural architecture search, or NAS), these quantities scale as:
@@
(a) Model Complexity vs Best Accuracy
(b) Model Complexity vs Accuracy Dist.
(c) Plot Legend
(d) Size Comparison between EffNet-B, EffNet-S, and LeNet
Figure 3: Performance comparison of different CNN architectures: fig. (a) shows the best accuracies vs. number of model parameters of the nets used, fig. (b) shows the median accuracies and error-bars in the same setting, fig. (c) is the legend for (a) and (b), and fig. (d) is a comparison of sizes of the networks compared to traditional EfficientNet versions.
where $\alpha\geq 1$, $\beta\geq 1$, $\gamma\geq 1$. Here, increasing $\phi$, the scaling parameter, increases the complexity of the network in a compounded manner. Noting that doubling the net-depth doubles the computational cost (floating point operations per second, or FLOPS), but doubling the other two quadruples the FLOPS, these also follow an empirical constraint: $\alpha\cdot\beta^{2}\cdot\gamma^{2}\simeq 2$, so that increasing $\phi$ by some $\phi_{0}$ increases the FLOPS by $2^{\phi_{0}}$. The authors of the first EfficientNet paper performed a grid search on $\alpha$, $\beta$, and $\gamma$ and obtained the values $1.2$, $1.1$, and $1.15$, respectively. Starting with $\phi=0$ and increasing $\phi$ by 1 each time, they obtained eight networks ($B_{0}$ to $B_{7}$) with increasing complexity. The base network ($B_{0}$) works on arrays of dimension $224\times 224$, while the largest of these models ($B_{7}$) takes in arrays of dimension $600\times 600$. There are seven main building blocks of an EfficientNet, each containing multiple parts. Building upon other successful lightweight models Sandler et al. (2018); Tan et al. (2019), EfficientNets replace a fully convolutional operation with a depthwise convolution (a single filter per channel) and a pointwise ($1\times 1$) convolution (see the blown up schematic of Block 2A in Fig. (2) ). This significantly reduces the computational cost while maintaining similar performance. When stacked together, the blocks mentioned above also perform a function known as inverted bottlenecking, embedding the input manifold into a lower-dimensional manifold. Each of these blocks also contains a squeeze and expansion block Hu et al. (2018), with a residual connection maintaining the information flow, as well as a combination of Swish, RELU, and Sigmoid activation functions. More details can be found in the original paper Tan and Le (2019). While the width coefficient scales up the channel dimension of each block, the depth coefficient determines the number of times each type of block is repeated, making the network deeper. These networks also scale up the dropout probability to regularize the larger models. All versions of EfficientNet outperform other CNNs that are much larger in size (in terms of accuracy) while being faster during inference ($B_{0}$, with 5.3M5.3~M parameters, achieves a top-5 accuracy of $93.3\%$ for ImageNet, compared to the widely used ResNet-50 (26M26~M parameters)—which has a top-5 accuracy of $93.0\%$; the FLOPS for ResNet-50 is $11$ times that of EfficientNet-$B_{0}$).
@@
Figure 4: ROC curve (signal efficiency vs. background rejection) for some of the networks (with some metric highest among others; table (2) has them in bold-face). The solid lines are for models working on both images and global features, dashed ones correspond to those with only images as input. The solid black curve is the optimal ROC curve obtainable by any model, obtained from likelihood ratio analysis (ref. Geuskens:2024tfo).
To circumvent this without changing the hyperparameters of the original EfficientNet architecture (and using the prescribed values of $\alpha$, $\beta$, and $\gamma$), we use negative $\phi$ values to scale down the input resolution. The smallest resolution thus achieved is $36\times 36$. Any smaller, and the bottlenecks within the network would not hold without modification. We choose the starting resolution to be $36\times 36$, corresponding to $\phi=-13$, then repeatedly increase $\phi$ by 1 to get the rest of the configurations. We thus create 5 variants, the last one corresponding to $\phi=-9$ [^6] Similar efforts can be found here: https://github.com/qubvel/efficientnet/issues/26. The other important parameter to set is the dropout probability, which is gradually increased from $B_{0}$ to $B_{7}$, as more regularization is needed for larger networks. We linearly extrapolated these values using the width and depth coefficients and found that for input array widths <90<90, we do not need dropouts. Altogether, we will denote this smaller architecture as EfficientNet-Small (EffNet-S) for this work. For these values of $\phi$, every block appears only once in the smaller EffNet-S structures. We have thus indexed the blocks ‘$i$-A’ ($i=1,\,2,\,\ldots$), to keep the possibility of larger networks open [^7] This raises an obvious question: as these CNNs were fine-tuned to detect specific shapes (with detail) from larger images, unlike in our case—where the jet-images are smaller and are mere representations with a much larger importance to finer structures—could it be possible to systematically search for and create versions of scalable networks for this specific task, albeit smaller? Though intriguing, this question is outside the scope of the present work. However, glimpses of this possibility may be perceived from the present analysis.. Table (1) lists the input dimensions for this architecture for two different input resolutions: $40\times 40$ and $64\times 64$.